# Revisiting, Benchmarking and Understanding Unsupervised Graph Domain Adaptation

### v1

@@ 5 Experimental Results and Analyses
In this section, we study the experimental results of all the models. We first provide a comprehensive comparison of specialized UGDA methods across five datasets with diverse distribution shifts. Following that, we perform a thorough analysis between different GNN variants to understand how data shift imposes challenges on GNNs. Finally, we present the performance of SimGDA variants, which shows the limit of GNNs’ transferability. For more information about metrics, hyperparameters, search spaces, and other implementation details, please refer to Appendix D.
5.1 Overall Comparisons
Observation 1: When facing significant shifts, it is important to design solutions tailored to mitigate structural discrepancies. Although several methods that incorporate classical DA approaches with deep node embedding have achieved impressive results on Airport, Blog and ArnetMiner datasets (e.g., ACDNE and DMGNN), they fail to obtain satisfied performance on datasets with significant shifts. These results underscore the limitations of marginal distribution alignment techniques in the presence of significant structural and label shifts in graph data. As shown in Table 3, methods tailored to address graph structure shifts show reasonable improvements over those incorporating classical DA methods with deep node embedding. This suggests that mitigating the impact of graph structure shift on node representation learning under this scenario is crucial.
@@ 5.2 Understanding and unlocking the inherent power of GNN
Note that despite many UGDA methods integrating traditional domain adaptation techniques, we observe their performance remains unsatisfactory and can even be inferior to that of SimGDA. This leads us to an intriguing question: do the intrinsic mechanisms of GNN play a more crucial role in enhancing transferability? To further investigate the question, we take one-layer GCN combined with vanilla DA as a baseline, and compare six variants: Max-Aggr, Mean-Aggr, GAT-Aggr, GIN-Aggr, with-no-neighbor, and with-multi-hop-neighbor; shown in shown in Figure 3. Moreover, we enhance above SimGDA with unsupervised graph learning techniques, referred to as SimGDA+, to explore the limit of GNNs in graph domain adaptation; shown in Table 2 and Table 3.
Observation 3: SimGDA achieves competitive performance compared with UGDA methods. As shown in Table 2 and Table 3, SimGDA shows comparable results with state-of-the-art approaches across various datasets. Specifically, SimGDA exhibits better performance than structure shift tailored solutions in datasets with minor data shifts, i.e. Airport, Blog and ArnetMinver, and demonstrates advantages surpassing state-of-the-art methods in E $\rightarrow$ U and U $\rightarrow$ E. Similar trends are also observed in the Twitch and MAG dataset, where SimGDA outperform most UGDA methods in six tasks. To further explore the impact of data shifts on GNNs, we analyze the role of aggregation scope and mechanisms, which tackle diverse shifts through the graph structure.
@@
Besides, we note a performance decline on the Airport and Blog datasets with the inclusion of multi-hop neighbors. Conversely, on the ArnetMiner, Twitch, and MAG datasets, incorporating multi-hop neighbors leads to performance improvement. The enhancement in performance in these cases can be attributed to the fact that heterophilous graphs exhibit a larger degree of conditional shift, and aggregation process may help to mitigate this situation by providing a more comprehensive view of the node’s context.
@@
To enhance the expressive power of aggregation operator, we evaluate aggregate mechanism used in GAT and GIN. Despite their stronger expressive capacities, their performance is generally ineffective across datasets such as ArnetMiner, Blog, Airport, and Twitch. This inefficiency can be attributed to the increased risk of model bias towards information from the source domain, stemming from the requirement of more parameters for learning. Consequently, source-bised discriminative aggregation mechanisms deteriorate the model’s transfer capability.
@@
Observation 6: GNNs can serve as a powerful graph domain adaptor. With the optimal number of neighbor hops and aggregators, we further enhance SimGDA with unsupervised graph learning techniques and obtain SimGDA+. As shown in Table 2 and Table 3, SimGDA+ surpasses the performance of most specially-designed methods for graph domain adaptation and even achieves the best performance in certain tasks. This superior performance under various types of shifts showcases the potential of GNNs as powerful domain adaptors. In summary, we contend that GNNs, when designed with appropriate aggregators, careful selection of neighbor hops, and the application of unsupervised graph learning techniques, can serve as effective and reliable graph domain adaptors.
@@ 6 Conclusion
In this paper, we introduce GDABench, the first comprehensive benchmark for unsupervised graph domain adaptation. Our evaluation encompasses 16 well-known models across 5 real-world datasets exhibiting diverse data distribution shifts. Furthermore, we design 6 GNN variants to investigate the inherent transferability of GNNs, enhancing them with 3 unsupervised techniques to explore their potential limits. Our empirical results shows that (1) the performance of current UGDA models varies significantly across different datasets and adaptation scenarios; (2) tailored strategies are essential for addressing and mitigating graph structural shifts, particularly when distribution discrepancies are substantial. (3) the transferability of GNNs in UGDA is heavily dependent on aggregation scope and architecture, influenced by factors such as label shift severity and graph heterophily. We also provide unified APIs and adopt consistent data processing as well as data splitting approaches for fair comparisons. In the future, we plan to extend GDABench to include broader scenarios [61], more cutting-edge models and more complex types of datasets. We hope our benchmark and findings will promote realistic and rigorous evaluations, inspiring new advances in graph domain adaptation.
@@ B.2 Shift Statistics of Datasets
According to dataset statistics, shown in Table 4 and Figure 4, we measure the degree of domain shift exhibited in the datasets for each tasks using statistical methods. We use MMD [40], CSS [25], Kullback-Leibler Divergence to characterize the degree of feature shift, structure shift and label shift. The results of each tasks is shown in Table 9. We take the average results of all tasks as the shift statistics for the datasets, shown in Table 5.
@@ D.1 Metrics
(a) Running time of B $\rightarrow$ U in Airport.
(b) Running time of D $\rightarrow$ A in ArnetMiner.
(c) Running time of FR $\rightarrow$ JP in MAG.
(d) Memory consumption of B $\rightarrow$ U in Airport.
(e) Memory consumption of D $\rightarrow$ A in ArnetMiner.
(f) Memory consumption of FR $\rightarrow$ JP in MAG.
Figure 6: Running time and memory consumption of baselines.
@@ D.2 Additional Experimental Details
Hardware Specifications.
The experiments are performed on one Linux server (CPU: Intel(R) Xeon(R) Platinum 8163 CPU @ 2.50GHz, Operation system: Ubuntu 18.04.5 LTS). For GPU resources, one NVIDIA Tesla V100 card with 32G are utilized The python libraries we use to implement our experiments are Python 3.8, Pytorch 1.13.1, Pytorch-geometric 2.4.0, Pytorch-sparse 0.6.15 and Pytorch-scatter 2.1.0.
Hyperparameter Settings.
To control the effect of hyperparameter selection and ensure fairness, we standardize the evaluation process with hyperparameter tuning. We utilize grid search to form the predefined search space for each models. We use all the source nodes and target nodes for model training. The experiments are repeated three times, and we report the mean performance. Table 10 provides a comprehensive list of all hyperparameters used in our grid search.
More Experimental Results
In accordance with Table 2 and 3, we provide the performance for all tasks of each model in Table 6, 7 and 8. In accordance with Figure 3, we provide the compared performance of vanilla DA with 6 GNN variants in Figure 5. We also demonstrate the running time and memory consumption of each model on S/M/L datasets respectively. For time consumption, we evaluate the efficiency of baselines by measuring the time it takes to converge. As shown in Figure 6, we can observe that some algorithms (e.g. A2GNN) can achieve relatively good performance with less complexity.

### v2

@@ 5 Experimental Results and Analyses
In this section, we study the experimental results of all the models. We first provide a comprehensive comparison of specialized UGDA methods across five datasets with diverse distribution shifts for node classification task. Following that, we perform a thorough analysis between different GNN variants to understand how data shift imposes challenges on GNNs. Finally, we present the performance of SimGDA variants, which shows the limit of GNNs’ transferability. For more information about metrics, hyperparameters, search spaces, and other implementation details, please refer to Appendix D. We further extend our scope to include the graph-level classification task; please refer to the Appendix E for details.
@@
5.1 Overall Comparisons
Observation 1: When facing significant shifts, it is important to design solutions tailored to mitigate structural discrepancies. Although several methods that incorporate classical DA approaches with deep node embedding have achieved impressive results on Airport, Blog and ArnetMiner datasets (e.g., ACDNE and DMGNN), they fail to obtain satisfied performance on datasets with significant shifts. These results underscore the limitations of marginal distribution alignment techniques in the presence of significant structural and label shifts in graph data. As shown in Table 3, methods tailored to address graph structure shifts show reasonable improvements over those that incorporate classical DA techniques with deep node embeddings. This suggests that mitigating the impact of graph structure shift on node representation learning under this scenario is crucial.
@@ 5.2 Understanding and unlocking the inherent power of GNN
Although many UGDA methods integrate traditional domain adaptation techniques, we observe their performance remains unsatisfactory and can even fall below that of SimGDA. This leads us to an intriguing question: do the intrinsic mechanisms of GNN play a more crucial role in enhancing transferability? To further investigate the question, we take one-layer GCN combined with vanilla DA as a baseline, and compare six variants: Max-Aggr, Mean-Aggr, GAT-Aggr, GIN-Aggr, with-no-neighbor, and with-multi-hop-neighbor; The results are shown in Figure 3. Moreover, we enhance above SimGDA with unsupervised graph learning techniques, referred to as SimGDA+, to explore the limit of GNNs in graph domain adaptation, which is shown in Table 2 and Table 3.
Observation 3: SimGDA achieves competitive performance compared with UGDA methods. As shown in Table 2 and Table 3, SimGDA shows comparable results with state-of-the-art approaches across various datasets. Specifically, SimGDA exhibits better performance than structure shift tailored solutions in datasets with minor data shifts, i.e., Airport, Blog and ArnetMinver, and demonstrates advantages surpassing state-of-the-art methods in E $\rightarrow$ U and U $\rightarrow$ E. Similar trends are also observed in the Twitch and MAG dataset, where SimGDA outperforms most UGDA methods in six tasks. To further explore the impacts of data shifts on GNNs, we analyze the role of aggregation scope and mechanisms, which tackle diverse shifts through the graph structure.
@@
Besides, we note a performance decline on the Airport and Blog datasets when including multi-hop neighbors. However, on the ArnetMiner, Twitch, and MAG datasets, incorporating multi-hop neighbors leads to performance improvement. The enhancement in performance in these cases can be attributed to the fact that heterophilous graphs exhibit a larger degree of conditional shift, and aggregation process may help to mitigate this situation by providing a more comprehensive view of the node’s context.
@@
To enhance the expressive power of aggregation operator, we evaluate aggregate mechanism used in GAT and GIN. Despite their stronger expressive capacities, their performance is generally ineffective across datasets such as ArnetMiner, Blog, Airport, and Twitch. This inefficiency can be attributed to the increased risk of model bias towards information from the source domain, stemming from the requirement of more parameters for learning. Consequently, source-biased discriminative aggregation mechanisms deteriorate the model’s transfer capability.
@@
Observation 6: GNNs can serve as a powerful graph domain adaptor. With the optimal number of neighbor hops and aggregators, we further enhance SimGDA with unsupervised graph learning techniques and obtain SimGDA+. As can be seen in Table 2 and Table 3, SimGDA+ surpasses the performance of most specially-designed methods for graph domain adaptation and even achieves the best performance in certain tasks. This superior performance under various types of shifts showcases the potential of GNNs as powerful domain adaptors. In summary, we contend that GNNs, when designed with appropriate aggregators, careful selection of neighbor hops, and the application of unsupervised graph learning techniques, can serve as effective and reliable graph domain adaptors.
5.3 Do LLMs help mitigate distribution shift in graphs?
Recently, Large Language Models (LLMs) [62] have demonstrated an impressive ability to understand and handle various text-related tasks. When dealing with text-attributed graph, an important question arises: does the distribution shift persist after leveraging LLMs as feature encoders?
To investigate this question, we utilize the prompts from TAPE [63], which allows us to assess the impact of LLM-based features on the model’s performance. For datasets, we choose the widely used ogbn-arxiv dataset [64], which contains paper title and abstract text information. Each node represents an arXiv paper, with directed edges indicating citations from one paper to another. The objective is to predict the 40 subject areas of arXiv computer science papers, which are manually assigned by the authors and moderators of arXiv. We split the data into three disjoint domains based on the publication year of the papers, i.e. 1950-2016, 2016-2018, 2018-2020. The statistical details for each domain are shown in the Table 4.
Table 4: Statistics of the split ogbn-arxiv dataset.
Domains ##\## Nodes ##\## Edges ##\## Homo ##\## Avg Degree
1950-2016 69,499 237,163 0.6945 3.41
2016-2018 51,241 111,754 0.6886 2.18
2018-2020 48,603 60,403 0.7092 1.24
We explored two approaches to enhance the original node attributes:
  • LLM enhanced text with word2vec embedding [65], which combines the title, abstract, and LLM-generated predictions and explanations into a single input. This composite text is then fed into word2vec. Then, the node features are obtained by averaging the embeddings of its combined input. We refer to it as arxiv-LLM-w2v.
  • LLM enhanced text with BERT embedding [66], which feeds the same composite text into a pretrained DeBERTa. Then, the node features are obtained by sentence embedding. Note that we did not finetune the DeBERTa like TAPE [63], since we focus on unsupervised graph domain adaptation. We refer to it as arxiv-LLM-bert.
First, we use MMD to characterize the degree of feature shift among these three datasets, i.e., ogbn-arxiv, arxiv-LLM-w2v and arxiv-LLM-bert. We consider 3 adaptation tasks, and the results are shown in Table 5. As we can see, when word2vec is used to encode the LLM-enhanced text, the feature shift is reduced compared to the original node features. However, when BERT is used for encoding, the feature shift increases. This indicates that the choice of text encoding method significantly influences the degree of feature shift.
Table 5: Feature shift among domains in arxiv dataset by using MMD [41] as metric.
Source Target ogbn-arxiv arxiv-LLM-w2v arxiv-LLM-bert
1950-2016 2016-2018 0.0405 $0$downarrow 0.0427 $\uparrow$
1950-2016 2018-2020 0.0528 0.0535 $\uparrow$ 0.0796 $\uparrow$
2016-2018 2018-2020 0.0148 0.0138 $\downarrow$ 0.0149 $\uparrow$
Next, we choose 5 recent graph domain adaptation models to assess the impact of LLM-based features on the model’s performance. Each experiment is repeated 3 times, and we report the average Micro-F1 score with standard deviation. As illustrated in Table 6, the performance of most baselines shows significant improvement with the arxiv-LLM-w2v dataset, whereas performance notably declines with the arxiv-LLM-bert dataset compared to the original ogbn-arxiv dataset. These results align with the MMD scores presented in the Table 5, which indicate that arxiv-LLM-bert exhibits a larger distribution shift compared to the other datasets. This emphasizes the necessity of selecting an appropriate text encoder when utilizing LLM-enhanced text for graph domain adaptation. An effective choice of text encoder can greatly impact the performance and mitigate the distribution shifts in text-attributed graph domain adaptation tasks.
Table 6: Micro-F1 score of 5 recent graph domain adaptation models with LLM-based features.
Source Target Features UDAGCN AdaGCN KBL GRADE A2GNN
1950-2016 2016-2018 ogbn-arxiv 52.42$\pm$0.52 61.78$\pm$0.18 50.88$\pm$0.24 61.85$\pm$0.17 60.18$\pm$0.54
arxiv-LLM-w2v 43.55$\pm$0.90 66.23$\pm$0.14 64.90$\pm$0.12 65.41$\pm$0.31 63.77$\pm$0.36
arxiv-LLM-bert 34.81$\pm$0.29 35.66$\pm$1.20 34.64$\pm$0.85 35.04$\pm$0.86 39.14$\pm$2.00
1950-2016 2018-2020 ogbn-arxiv 48.41$\pm$0.64 56.74$\pm$0.34 47.53$\pm$1.15 57.19$\pm$0.26 58.89$\pm$0.28
arxiv-LLM-w2v 39.56$\pm$0.54 62.04$\pm$0.29 60.67$\pm$0.29 62.04$\pm$0.29 65.35$\pm$0.12
arxiv-LLM-bert 30.18$\pm$0.25 31.69$\pm$0.42 30.26$\pm$1.09 31.09$\pm$0.12 35.40$\pm$1.55
2016-2018 2018-2020 ogbn-arxiv 54.84$\pm$0.22 62.05$\pm$0.04 52.99$\pm$0.05 61.42$\pm$0.14 59.45$\pm$0.28
arxiv-LLM-w2v 47.20$\pm$0.99 68.42$\pm$0.04 66.77$\pm$0.13 66.84$\pm$0.16 65.83$\pm$0.40
arxiv-LLM-bert 39.20$\pm$2.53 39.93$\pm$1.90 37.27$\pm$2.06 34.14$\pm$1.54 43.62$\pm$1.87
@@ 6 Conclusion
In this paper, we introduce GDABench, the first comprehensive benchmark for unsupervised graph domain adaptation. Our evaluation encompasses 16 well-known models across various real-world datasets exhibiting diverse data distribution shifts. Furthermore, we also designed 6 GNN variants to investigate the inherent transferability of GNNs, enhancing them with 3 unsupervised techniques to explore their potential limits. Our empirical results shows that (1) the performance of current UGDA models varies significantly across different datasets and adaptation scenarios; (2) tailored strategies are essential for addressing and mitigating graph structural shifts, particularly when distribution discrepancies are substantial. (3) the transferability of GNNs in UGDA is heavily dependent on aggregation scope and architecture, influenced by factors such as label shift severity and graph heterophily. We have provided unified APIs and adopted consistent data processing as well as data splitting approaches for fair comparisons. In the future, we plan to extend GDABench to include broader scenarios [67, 68], more cutting-edge models and more complex types of datasets. We hope our benchmark and findings will promote realistic and rigorous evaluations, inspiring new advances in graph domain adaptation.
@@
Acknowledgments and Disclosure of Funding
This work is supported by the National Natural Science Foundation of China (62372408, 62106221), Zhejiang Provincial Natural Science Foundation of China (Grant No: LTGG23F030005), Ningbo Natural Science Foundation (Grant No: 2022J183). The research of Zhen Zhang and Bingsheng He is supported by the National Research Foundation, Singapore under its Industry Alignment Fund – Pre-positioning (IAF-PP) Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Section 1.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Section 6.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] Our paper strictly conforms to the ethics review guidelines.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] Our paper does not include theoretical results.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] Our paper does not include theoretical results.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See URL in Abstract.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Appendix D.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Table 2, Table 3 and Appendix D.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix D.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See footnotes and references.
    2. (b)
      Did you mention the license of the assets? [Yes] We use MIT license for our released assets.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] Our codes and datasets are available through URLs.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Appendix B.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We use public datasets that do not contain personally identifiable information or offensive content.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] Our research do not involve human subjects.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] Our research do not involve human subjects.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] Our research do not involve human subjects.
@@ B.2 Shift Statistics of Datasets
According to dataset statistics, shown in Table 7 and Figure 4, we measure the degree of domain shift exhibited in the datasets for each tasks using statistical methods. We use MMD [41], CSS [26], Kullback-Leibler Divergence to characterize the degree of feature shift, structure shift and label shift. The results of each tasks is shown in Table 13. We take the average results of all tasks as the shift statistics for the datasets, shown in Table 8. The 74 tasks compiled by the five carefully selected datasets can cover all combinations of domain shift scenarios.
[…]
@@ D.1 Metrics
(a) Group 1
(b) Group 2
(c) Group 3
(d) Group 1
(e) Group 2
(f) Group 3
[…]
@@ D.2 Additional Experimental Details
  • Hardware Specifications. The experiments were conducted on a Linux server equipped with an Intel(R) Xeon(R) Platinum 8163 CPU operating at 2.50GHz, running Ubuntu 18.04.5 LTS. For GPU resources, we utilized a single NVIDIA Tesla V100 graphics card with 32GB of memory. The Python libraries employed for implementing our experiments include Python 3.8, PyTorch 1.13.1, PyTorch Geometric 2.4.0, PyTorch Sparse 0.6.15, and PyTorch Scatter 2.1.0.
  • Hyperparameter Settings. To control the effect of hyperparameter selection and ensure fairness, we standardize the evaluation process with hyperparameter tuning. We utilize grid search to form the predefined search space for each models. We use all the source nodes and target nodes for model training. The experiments are repeated three times, and we report the mean performance. Table 14 provides a comprehensive list of all hyperparameters used in our grid search.
  • More Experimental Results. In accordance with Table 2 and 3, we provide the performance for all tasks of each model in Table 9, 11 and 12. In accordance with Figure 3, we provide the compared performance of vanilla DA with 6 GNN variants in Figure 5.
  • Exploration of Hyperparameter Impact. We investigate how various hyperparameters in common modules influence the performance of different UGDA methods on ArnetMiner dataset (task D $\rightarrow$ A). We focus on two key aspects: the number of GNN layers and the representation dimensions. Results are shown in Figure 6.
  • Running Time and Memory Consumption. We also demonstrate the running time and memory consumption of each model on S/M/L datasets respectively. For time consumption, we evaluate the efficiency of baselines by measuring the time it takes to converge. As shown in Figure 7, we can observe that some algorithms (e.g. A2GNN) can achieve relatively good performance with less complexity.
@@ D.3 The PyGDA Library
Appendix E Experiments on Graph Classification
To expand our research scope, we take graph-level shifts into consideration and add a pooling layer to evaluate capabilities of baselines in graph-level domain adaptation. We employ three TUdatasets: Proteins, Mutagenicity, and Frankenstein, partitioning each dataset into 2 equally sized disjoint groups based on density shifts. Detailed statistics are shown in Table 10.
Table 10: Statistics of graph-level datasets in GDABench.
Dataset ##\## Nodes ##\## Edges ##\## Feature ##\## Class Num of graphs
Proteins  39.06  72.82 4 2 1,113
Mutagenicity  30.32  30.77 14 2 4,337
Frankenstein  16.90  17.88 780 2 4,337
(a) Running time of B $\rightarrow$ U in Airport.
(b) Running time of D $\rightarrow$ A in ArnetMiner.
(c) Running time of FR $\rightarrow$ JP in MAG.
(d) Memory consumption of B $\rightarrow$ U in Airport.
(e) Memory consumption of D $\rightarrow$ A in ArnetMiner.
(f) Memory consumption of FR $\rightarrow$ JP in MAG.
Figure 7: Running time and memory consumption of baselines.
@@ Appendix E Experiments on Graph Classification
The results are detailed in Table 15 and Table 16. Among the methods, GRADE and A2GNN are domain adaptive message passing methods and the remaining are DA incorporated node embedding methods. Key observations are as follows:
DA incorporated node embedding methods shows task-inconsisteny across node and graph-level tasks. For example, DANE performs averagely in node-level tasks, but its performance improves significantly in graph-level tasks. This disparity highlights a challenge in predicting the performance of unsupervised graph domain adaptation (UGDA) models in real-world applications. The inconsistency suggests that models optimized for node-level tasks may not generalize well to graph-level tasks and vice versa. Consequently, this variability complicates the task of assessing how well these models will perform when deployed in diverse and complex real-world scenarios where both node-level and graph-level information may be critical.
Domain adaptive message passing methods demonstrate superior and consistency performance across a wide range of datasets and tasks. As shown in Table 3, 9, 16 and 15, methods designed based on the inherent properties of GNN achieves the top-three best performance in 8 tasks out of 12 node-level tasks and top-two best performance in 5 tasks out of 6 graph-level tasks. This observation verified our findings that establishing domain adaptation principles by leveraging inherent properties of GNN can result in an effective and efficient approach to addressing the challenges of domain variability in graph datasets.
To summarize, our observations underscore the importance of leveraging the intrinsic properties of GNNs to devise effective domain adaptation strategies, which not only enhances performance but also ensures consistency in real-world applications.
Table 15: To evaluate the baselines on graph-level shifts, we compared the Micro-F1 scores of each model on the Proteins, Mutagenicity, and Frankenstein datasets. The best results are highlighted in bold, and the second-best results are underlined.
Proteins Mutagenicity Frankenstein
Models P1 $\rightarrow$ P2 P2 $\rightarrow$ P1 M1 $\rightarrow$ M2 M2 $\rightarrow$ M1 F1 $\rightarrow$ F2 F2 $\rightarrow$ F1
DANE 60.14 ±3.58 75.66 ±0.98 67.25 ±0.14 76.92 ±0.35 54.77 ±0.53 56.96 ±2.89
UDAGCN 53.50 ±2.42 73.14 ±4.29 58.11 ±0.58 65.34 ±0.55 52.48 ±0.32 52.37 ±1.38
AdaGCN 52.60 ±0.78 78.12 ±0.37 58.89 ±0.06 56.18 ±0.02 56.28 ±0.75 53.01 ±3.63
CWGCN 50.45 ±4.81 44.84 ±8.20 55.60 ±1.27 56.72 ±0.67 49.76 ±0.27 51.92 ±0.71
SAGDA 53.14 ±4.80 46.22 ±2.99 57.06 ±3.54 56.00 ±8.85 50.35 ±0.26 51.01 ±8.37
GRADE 43.93 ±0.31 76.80 ±0.29 69.00 ±0.22 76.57 ±0.31 57.54 ±1.09 58.39 ±4.57
A2GNN 51.70 ±1.54 69.65 ±4.21 56.83 ±0.19 58.88 ±1.23 50.43 ±0.69 48.99 ±3.97
Table 16: To evaluate the baselines on graph-level shifts, we compared the Macro-F1 scores of each model on the Proteins, Mutagenicity, and Frankenstein datasets. The best results are highlighted in bold, and the second-best results are underlined.
Proteins Mutagenicity Frankenstein
Models P1 $\rightarrow$ P2 P2 $\rightarrow$ P1 M1 $\rightarrow$ M2 M2 $\rightarrow$ M1 F1 $\rightarrow$ F2 F2 $\rightarrow$ F1
DANE 59.14 ±3.06 56.30 ±6.09 67.11 ±0.17 76.50 ±0.35 52.24 ±1.02 54.94 ±2.13
UDAGCN 53.15 ±2.74 50.19 ±1.20 56.71 ±0.61 63.35 ±0.56 50.06 ±0.64 52.32 ±1.40
AdaGCN 49.33 ±1.62 57.99 ±2.82 58.00 ±0.10 35.97 ±0.10 55.99 ±0.94 51.76 ±4.43
CWGCN 40.57 ±3.13 42.75 ±6.01 39.00 ±4.96 37.32 ±1.66 39.46 ±0.22 51.68 ±0.67
SAGDA 46.65 ±6.14 33.42 ±1.01 56.26 ±3.74 54.95 ±8.22 36.89 ±4.93 38.03 ±6.81
GRADE 32.23 ±0.86 50.52 ±1.77 68.98 ±0.21 76.32 ±0.26 56.93 ±1.80 54.98 ±2.62
A2GNN 47.71 ±3.22 58.85 ±1.16 55.42 ±0.10 50.17 ±1.59 46.97 ±0.96 43.33 ±1.87
Appendix F Discussion
F.1 How these findings generalize to real-world scenarios
Our benchmark includes a range of datasets with varying characteristics to capture different aspects of graph domain adaptation. This diversity aims to provide a broad perspective on the applicability of our methods. In real-world scenarios, applying graph adaptation methods effectively involves several key considerations: Firstly, it is imperative to develop tailored strategies specifically designed to address the structural shifts observed in graphs. For example, if a graph is dynamic and changing overtime, it is crucial to accord greater attention to its evolving structure. Secondly, recognizing the importance of the aggregation scope and aggregation architecture in GNNs’ transferability within unsupervised graph domain adaptation (UGDA) are crucial. In real-world graphs, noise is inevitable, hence, strategically selecting effective neighbors not only improve performance but also avoid noise. Thirdly, by leveraging the properties of GNNs that make them inherently adaptable to changes in graph structure and data distribution, we can develop simple yet highly effective models.
F.2 A broader discussion on DA problem and other related UGDA scenarios
UDA vs UGDA.
Unsupervised domain adaptation (UDA) entails transferring knowledge from a labeled source domain to an unlabeled target domain. A prevalent strategy in domain adaptation is to reduce domain discrepancies while learning domain-invariant representations, a method that has seen considerable success in the fields of computer vision and natural language processing. However, these techniques typically operate under the assumption that inputs are independently and identically distributed (IID), making them unsuitable for tasks involving non-IID data, such as node classification in graph-structured datasets.
UGDA vs muti-domain UGDA.
Muti-domain UGDA extends the concept of domain adaptation to situations where there are multiple source domains and a single target domain. This approach aims to learn a model that can generalize well across multiple source domains, and then adapt it to perform well on the target domain. Compared to standard UGDA, multi-domain UGDA can enhance generalization by leveraging the diversity of multiple source domains. However, it may require more complex models and additional computational resources.
UGDA vs source-free UGDA.
Source-free UGDA advances domain adaptation by tackling the challenge of adapting models without access to labeled data from the source domains. This setting is more challenging as it involves learning to transfer knowledge without explicit supervision. Source-free UGDA methods often employ techniques such as self-training or consistency regularization to adapt the model to the target domain. Compared to UGDA, source-free UGDA may be more sensitive to domain shift and require careful selection of adaptation techniques.